OpenStudy (anonymous):

Solve the following system using substitution or elimination: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Be sure to show all work.

OpenStudy (anonymous):

y = x + 4 x + x + x + 4 = 46 3x + 4 = 46 3x = 42 x = 14 y = x + 4 = 18 Sides are 14, 14, and 18